Circular No. 6098

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

λ ERIDANI

S. Stefl, Astronomical Institute, Czech Academy of Sciences, Ondřejov, communicates: “Two Reticon observations (620.0–672.0 nm) of the wellknown B2e IV star λ Eri were obtained at the coudé focus of the Ondřejov 2-m telescope on Oct. 19 and 21. They show extraordinarily strong emission in Hα, He I (667.8 nm) and Si II. Hα has a double-peaked emission profile with peak intensities of 1.37 and 1.38, respectively, of the continuum level. The bottom of the shell absorption declines to about half that of the emission, corresponds to a radial velocity of 12 km/s and shows different asymmetry in two spectra separated by two days. The violet and red emission wings of He I (667.8 nm) rise 2 percent above the continuum level. Emission up to 0.5 percent of the continuum can be recognized also in the Si II 634.3- and 637.1-nm lines. The last emission activity of the star was observed in Jan. 1994 (Stefl 1994, Be Star Newsl. No. 28, p. 5). A concentrated observing effort is desirable, assuming that the new emission activity persists — like the previous one — only for up to a few months. A multi-site observing campaign on λ Eri is scheduled in the period 1994 Nov. 12–18 (Hirata 1994, ibid. No. 28, p. 11).”

SATURN J. Lancashire reports his timings (± 0.003 day) of the transit of the

white equatorial spot (IAUC 6079, 6092) on Saturn with the 0.30-m refractor at the University of Cambridge Observatories, and he notes that these are consistent with a rotation rate of 10^{h}22^{m} for the spot: p (preceding edge), Oct. 13.842 UT; c (center), 13.851; f (following edge), 13.860; p, 22.915; c, 22.922; f, 22.932; c, 23.788; f, 13.800; c, 25.944; f, 25.953. The spot has appeared as bright as the brightest part of the rings, with the rest of the equatorial zone slightly shaded and the equatorial belts darker than usual. The spot has shown no significant change over the past two dozen or so rotations.

B. L. Greiner, East Petersburg, PA, reports his timings for transits of this white spot: Oct. 16.010 UT; 18.170; 19.035.

POSSIBLE GRAVITATIONAL MICROLENSING EVENT Further to his item on IAUC 6097, E. Giraud adds: “The luminosity of

the microlens candidate seems to have reached a maximum. The magnitude difference between Oct. 18–19 and 20–21 is only 0.04 ± 0.08 mag.”

1994 October 26 Gareth V. Williams